# Nesterov’s Accelerated Jacobi-Type Methods for Large-scale Symmetric Positive Semidefinite Linear Systems

### v1

@@ Abstract
Solving symmetric positive semidefinite linear systems is an essential task in many scientific computing problems. While Jacobi-type methods, including the classical Jacobi method and the weighted Jacobi method, exhibit simplicity in their forms and friendliness to parallelization, they are not attractive either because of the potential convergence failure or their slow convergence rate. This paper aims to showcase the possibility of improving classical Jacobi-type methods by employing Nesterov’s acceleration technique that results in an accelerated Jacobi-type method with improved convergence properties. Simultaneously, it preserves the appealing features for parallel implementation. In particular, we show that the proposed method has an $O\left(\frac{1}{t^{2}}\right)$ convergence rate in terms of objective function values of the associated convex quadratic optimization problem, where $t\geq 1$ denotes the iteration counter. To further improve the practical performance of the proposed method, we also develop and analyze a restarted variant of the method, which is shown to have an $O\left(\frac{(\log_{2}(t))^{2}}{t^{2}}\right)$ convergence rate when the coefficient matrix is positive definite. Furthermore, we conduct appropriate numerical experiments to evaluate the efficiency of the proposed method. Our numerical results demonstrate that the proposed method outperforms the classical Jacobi-type methods and the conjugate gradient method and shows a comparable performance as the preconditioned conjugate gradient method with a diagonal preconditioner. Finally, we develop a parallel implementation and conduct speed-up tests on some large-scale systems. Our results indicate that the proposed framework is highly scalable.
@@ 1 Introduction
Solving SPSD linear systems is ubiquitous for obvious reasons in various fields of science and engineering. Many physical and engineering problems are naturally described by mathematical models involving SPSD matrices. For example, the Laplacian matrix of an undirected graph is SPSD, and it provides a highly effective mathematical framework for analyzing the structure and behavior of networks, graphs, and systems in a wide range of disciplines [9, 29, 37, 32, 31]. Another representative problems is to solve and analyze an elliptic partial differential equation (PDE). The stiffness matrices associated with those PDEs from finite element analysis [20], solid and structural mechanics [46], electrostatics [35, 41], and heat and mass transfer problems [17], are typically symmetric positive definite (SPD). Solving linear systems arising from these systems efficiently and reliably is crucial for exploring the behavior of these PDEs and making reliable predictions. Furthermore, SPSD/SPD matrices frequently arise in optimization problems and machine learning tasks, both of which are foundational tools for addressing a wide range of real-world applications. In particular, algorithmic frameworks for solving constrained and unconstrained optimization problems [6] often involve solving SPSD/SPD linear systems at their core. Efficient solution methods for such systems are essential for successful applications.
@@
For small to medium-sized problems, one can generally apply a direct method [13], including Cholesky factorization and LU decomposition methods, to obtain solutions with high precision. Moreover, it is generally robust even if the matrix is ill-conditioned or nearly singular. However, as the computational complexity and memory consumption of a direct method are typically $O(n^{3})$ and $O(n^{2})$, respectively [^1] The complexity could be smaller when the matrix $Q$ is sparse but fill-ins would be introduced in general., it can become prohibitively expensive for solving large-scale linear systems. On the other hand, iterative methods [34, 14] require much smaller per-iteration computational complexity and are much more memory efficient, which make them particularly suitable for solving large-scale and/or sparse systems. Some notable iterative methods include Krylov subspace methods, classical Jacobi method, Gauss-Seidel method, (symmetric) successive over-relaxation method, Kaczmarz method, and their variants [13, 19, 44, 14, 21, 38], to mention just a few. Iterative methods can also be terminated to achieve different levels of accuracy, which is an advantage when highly accurate solutions are not necessary. However, the convergence of iterative methods is sensitive to the spectral property (and hence conditioning) of the coefficient matrix, and they generally require finding effective preconditioners to speed up the convergence, especially when the matrix is ill-conditioned. Furthermore, they may require too many iterations or even fail to reach a solution with a desired precision under the influence of rounding errors. But despite the shortcomings, iterative methods remain a powerful tool, and sometimes the only tool, for solving large-scale linear systems.
@@
In this paper, we focus on Jacobi-type frameworks in order to take advantage of their simple forms and the ease of parallelization. It is well-known that the classical Jacobi method does not converge for general SPSD/SPD systems. It is only guaranteed to be convergence under strong conditions, such as strict diagonal dominance of $Q$ [13, 34]. A modification of the classical Jacobi method by considering a linear combination of the Jacobi update and the current iterate leads to the weighted Jacobi method. The latter method admits convergence for any SPD system, but the weight needs to be chosen carefully [34]. As we will also see in Section 2, both Jacobi and weighted Jacobi methods can be viewed as special cases of the variable-metric gradient method for solving the associated convex QP problem, which requires suitable conditions for the selection of ”step sizes” for convergence. However, both methods typically show slow convergence rates in practice (when they are convergent).
We propose to accelerate the Jacobi-type methods by leveraging Nesterov’s acceleration technique [27, 3, 26, 39] together with a carefully designed proximal term. The resulting algorithm has a similar updating rule as Jacobi-type methods, which allows efficient and simple parallel implementation, and it is provably convergent under mild conditions. In particular, the proposed method is shown to be convergence with an $O\left(\frac{1}{t^{2}}\right)$ convergence rate in terms of objective function values of the associated convex QP problem, where $t\geq 1$ denotes the iteration counter. Moreover, existing literature also suggests that a restarted scheme could improve the practical performance of the proposed method significantly. Hence, we also analyze the convergence properties of a restarted version of the algorithm and establish an $O\left(\frac{(\log_{2}t)^{2}}{t^{2}}\right)$ convergence rate. Finally, our numerical experiments confirm the efficiency of the proposed algorithm and the scalability of its parallel implementation. Specifically, when applied to strictly diagonal dominant systems, our accelerated algorithm empirically outperforms both the classical Jacobi method and weighted Jacobi method with optimal weight. Furthermore, in numerical tests involving Laplacian systems and general SPD systems, our method demonstrates better empirical performance compared to the classical CG method, and shows comparable results as the preconditioned CG method using a simple diagonal preconditioner. Additionally, the speed-up tests for the parallel implementation of our method exhibit promising scalability. These combined theoretical and numerical results provide compelling evidence for the potential applicability of our algorithm in a wide range of real-world applications.
@@ 2 The Nesterov’s accelerated method
In this section, we first revisit the Jacobi-type algorithmic frameworks and show that they are special cases of the general variable-metric gradient methods for solving the associated unconstrained convex QP problem. Hence, the convergence results from the optimization perspectives can be directly applied. Then, we consider the Nesterov’s accelerated gradient method with provable convergence. We also present the convergence results of a restarted version of the former algorithm. We should mention that the analysis for the Nesterov’s accelerated gradient method and its restarted variants is well-developed and mature in the literature; see e.g., [5, 3, 27, 18, 24, 30, 43, 23] and references therein. However, existing analysis is mainly for a general convex optimization problem, whereas the focus of the present paper is an unconstrained convex QP problem. To make the paper self-contained, we also provide a concise analysis here.
@@
where $t$ is the iteration counter. The standard condition based on the fixed-point theory for the Jacobi method to be convergent is $\rho(D^{-1}(L+L^{T}))<1$ where $\rho(\cdot)$ denotes the spectral radius [13]. As a consequence, the Jacobi method is not always convergent for general SPD/SPSD systems given an arbitrary initial point, as mentioned before. In order to derive convergent Jacobi-type methods, one may consider a linear combination between the current point $x^{t}$ and the point given by the Jacobi method:
@@
where $\omega\in\left(0,\frac{2}{\lambda_{\textrm{max}}(D^{-1}Q)}\right)$. The above scheme is usually called the weighted Jacobi method, and it is convergent for any given initial point [34] when solving SPD systems. We can derive the optimal weight, denoted as $\omega_{\textrm{opt}}$ by minimizing the spectral radius $\rho(I-\omega D^{-1}Q)$ and it is given by
@@
(5) $x^{t+1}=\operatorname{argmin}\left\{f(x^{t})+\left\langle\nabla f(x^{t}),x-x^{t}\right\rangle+\frac{1}{2}\left\lVert x-x^{k}\right\rVert_{S}^{2}\;:\;x\in\mathbb{R}^{n}\right\},\quad\forall\;t\geq 0,$
where $S:=D$ for (3) and $S:=\frac{1}{\omega}D$ for (4), provided that $D\in\mathbb{S}_{++}^{n}$ and $\omega>0$. Here, the term $\frac{1}{2}\left\lVert x-x^{t}\right\rVert_{S}^{2}$ is usually called the proximal term. Note that the traditional gradient method corresponds to the setting $S=\frac{1}{\eta}I_{n}$, where $\eta>0$ is the step size. Therefore, the above updating rule is a special case of the gradient method by changing the standard Euclidean metric to the metric induced by $S$. Hence, these types of methods are also known as variable-metric gradient methods. However, the gradient method specified by (5) is not always convergence. In other words, one needs to choose the matrix $S$ appropriately for convergence. For a given $S\in\mathbb{S}_{++}^{n}$, observe that the updating rule in (5) has the following form:
@@
where $x^{*}$ is the unique solution of (1). The above convergence results coincide with those classical results in the field of optimization; see, e.g., [28, Theorem 2.1.15]. Unfortunately, when the condition number of $Q$, namely $\frac{\lambda_{\textrm{max}}(Q)}{\lambda_{\textrm{min}}(Q)}$, is large, the convergence rate could be rather slow. Moreover, when $Q$ is only SPSD, the above results are no longer applicable. In this case, we can still rely on the convergence results for convex optimization. For instance, suppose that $Q\in\mathbb{S}_{+}^{n}$ and $S=\frac{1}{\eta}I_{n}$, and $\eta=\frac{1}{\lambda_{\textrm{max}}(Q)}$, then
@@
which establishes an $O\left(\frac{1}{t}\right)$ convergence rate in terms of the objective value gap; see [28, Corollary 2.1.2] for more details. Note that the step size $\eta$ depends on $\lambda_{\textrm{max}}(Q)$ which requires additional computational cost to estimate.
@@
The Nesterov’s accelerated framework is recognized to be one of the most influential and efficient first-order methods in the literature. By the description of Algorithm 1, we see that the algorithm is easily applicable and simple to implement. Moreover, there are several major differences between Algorithm 1 and the gradient method (5). First, instead of using a linear approximation, i.e., $f(x^{t})+\left\langle\nabla f(x^{t}),x-x^{t}\right\rangle$, of the objective function $f$ at the current iterate $x^{t}$, we keep using the function $f$ that is in fact an exact second-order approximation. Second, the matrix $S$ is no longer required to be SPD. Instead, only $S\in\mathbb{S}_{+}^{n}$ is required. Finally, the proximal term is changed to $\frac{1}{2}\left\lVert x-y^{t}\right\rVert_{S}^{2}$ where $y^{t}$ depends on $x^{t-1}$ and $x^{t-2}$ which essentially adds a momentum term to the update of the gradient method (5) to achieve acceleration. The above features are some of the main reasons to achieve accelerated convergence rate, as we will see shortly.
@@ 2.1 Convergence properties
$f(x)-f(x^{t})\geq$ $\;f(x)-q(x^{t},y^{t})$
$=$ $\;\frac{1}{2}\left\langle x,Qx\right\rangle-\left\langle b,x\right\rangle-\frac{1}{2}\left\langle x^{t},Qx^{t}\right\rangle+\left\langle b,x^{t}\right\rangle-\frac{1}{2}\left\langle x^{t}-y^{t},S(x^{t}-y^{t})\right\rangle$
@@
where we have used the facts that $f(x)-f^{*}\geq 0$ for $x\in\mathbb{R}^{n}$ and that $\alpha_{t-1}^{2}=\alpha_{t}(\alpha_{t}-1)$, for $t\geq 1$. Then, by multiplying $\alpha_{t}$ on both side of the last inequality, we obtain that
@@
$=$ $\;\frac{1}{2}\left\langle\alpha_{t}x^{t}-(\alpha_{t}-1)x^{t-1}-x^{*},S(\alpha_{t}x^{t}-(\alpha_{t}-1)x^{t-1}-x^{*})\right\rangle$
$\;-\left\langle\alpha_{t}y^{t}-(\alpha_{t}-1)x^{t-1}-x^{*},S(\alpha_{t}y^{t}-(\alpha_{t}-1)x^{t-1}-x^{*})\right\rangle$
@@
where in the last equality we used the fact that $\alpha_{t}y^{t}-(\alpha_{t}-1)x^{t-1}-x^{*}=w^{t-1}$, for $t\geq 1$, which can be deduced directly from Line 4 and Line 5 in Algorithm 1.
@@
$0\leq f(x^{t})-f^{*}\leq\frac{2\left\lVert x^{0}-x^{*}\right\rVert_{S}^{2}}{(t+1)^{2}},\quad t\geq 1$
@@
The statement can be obtained easily by using the facts $\alpha_{0}=0$, $w^{0}=x^{0}-x^{*}$ and that $\alpha_{t}\geq\frac{t+1}{2}$, for all $t\geq 1$. Note that the latter fact can be proven by induction and using the updating rule of $\alpha_{t}$ as given in Line 4 of Algorithm 1. Hence, the proof is completed.
Theorem 2.5 gives an improved $O\left(\frac{1}{t^{2}}\right)$ convergence rate that is obviously better than the traditional gradient method, whose worst-case convergence rate is $O\left(\frac{1}{t}\right)$. However, careful readers may observe that the generated objective values $\{f(x^{t})\}$ are not necessarily monotonically decreasing. This is reasonable since Theorem 2.5 only provides a sequence of upper bounds for the objective values.

### v2

@@ Abstract
Solving symmetric positive semidefinite linear systems is an essential task in many scientific computing problems. While Jacobi-type methods, including the classical Jacobi method and the weighted Jacobi method, exhibit simplicity in their forms and friendliness to parallelization, they are not attractive either because of the potential convergence failure or their slow convergence rate. This paper aims to showcase the possibility of improving classical Jacobi-type methods by employing Nesterov’s acceleration technique that results in an accelerated Jacobi-type method with improved convergence properties. Simultaneously, it preserves the appealing features for parallel implementation. In particular, we show that the proposed method has an $O\left(t^{-2}\right)$ convergence rate in terms of objective function values of the associated convex quadratic optimization problem, where $t\geq 1$ denotes the iteration counter. To further improve the practical performance of the proposed method, we also develop and analyze a restarted variant of the method, which is shown to have an $O\left((\log_{2}(t))^{2}t^{-2}\right)$ convergence rate when the coefficient matrix is positive definite. Furthermore, we conduct appropriate numerical experiments to evaluate the efficiency of the proposed method. Our numerical results demonstrate that the proposed method outperforms the classical Jacobi-type methods and the conjugate gradient method, and shows a comparable performance as the preconditioned conjugate gradient method with a diagonal preconditioner. Finally, we develop a parallel implementation and conduct speed-up tests on some large-scale systems. Our results indicate that the proposed framework is highly scalable.
@@ 1 Introduction
Solving SPSD linear systems is ubiquitous for obvious reasons in various fields of science and engineering. Many physical and engineering problems are naturally described by mathematical models involving SPSD matrices. For example, the Laplacian matrix of an undirected graph is SPSD, and it provides a highly effective mathematical framework for analyzing the structure and behavior of networks, graphs, and systems in a wide range of disciplines [10, 33, 41, 36, 35]. Another representative problems is to solve and analyze an elliptic partial differential equation (PDE). The stiffness matrices associated with those PDEs from finite element analysis [23], solid and structural mechanics [51], electrostatics [39, 45], and heat and mass transfer problems [19], are typically SPSD. Solving linear systems arising from these systems efficiently and reliably is crucial for exploring the behavior of these PDEs and making reliable predictions. Furthermore, SPSD matrices frequently arise in optimization problems and machine learning tasks, both of which are foundational tools for addressing a wide range of real-world applications [6, 17, 22]. In particular, algorithmic frameworks for solving constrained and unconstrained optimization problems [6, 17] often involve solving SPSD linear systems at their core. Efficient solution methods for such systems are essential for successful applications.
@@
For small to medium-sized problems, one can generally apply a direct method [14], including Cholesky factorization and LU decomposition methods, to obtain solutions with high precision. Moreover, it is generally robust even if the matrix is ill-conditioned or nearly singular. However, as the computational complexity and memory consumption of a direct method are typically $O(n^{3})$ and $O(n^{2})$, respectively [^1] The complexity could be smaller when the matrix $Q$ is sparse but fill-ins would be introduced in general., it can become prohibitively expensive for solving large-scale linear systems. On the other hand, iterative methods [38, 15] require much smaller per-iteration computational complexity and are much more memory efficient, which make them particularly suitable for solving large-scale and/or sparse systems. Some notable iterative methods include Krylov subspace methods, classical Jacobi method, Gauss-Seidel method, (symmetric) successive over-relaxation method, Kaczmarz method, and their variants [14, 21, 49, 15, 24, 42], to mention just a few. Iterative methods can also be terminated to achieve different levels of accuracy, which is an advantage when highly accurate solutions are not necessary. However, the convergence of iterative methods is sensitive to the spectral property (and hence the conditioning) of the coefficient matrix, and they generally require finding effective preconditioners to speed up the convergence, especially when the matrix is ill-conditioned. Furthermore, they may require too many iterations or even fail to reach a solution with a desired precision under the influence of rounding errors. But despite the shortcomings, iterative methods remain a powerful tool, and sometimes the only tool, for solving large-scale linear systems.
@@
In this paper, we focus on Jacobi-type frameworks in order to take advantage of their simple forms and the ease of parallelization. It is well-known that the classical Jacobi method does not converge for general SPSD systems. It is only guaranteed to be convergent under strong conditions, such as strict diagonal dominance of $Q$ [14, 38]. A modification of the classical Jacobi method by considering a linear combination of the Jacobi update and the current iterate leads to the weighted Jacobi method. The latter method admits convergence for any symmetric positive definite system, but the weight needs to be chosen carefully [38]. As we will also see in Section 2, both Jacobi and weighted Jacobi methods can be viewed as special cases of the variable-metric gradient method for solving the associated convex QP problem, which requires suitable conditions for the selection of ”step sizes” for convergence. However, both methods typically show slow convergence rates in practice (when they are convergent).
We propose to accelerate the Jacobi-type methods by leveraging Nesterov’s acceleration technique [31, 3, 30, 43] together with a carefully designed proximal term. The resulting algorithm has a similar updating rule as Jacobi-type methods, which allows efficient and simple parallel implementation, and it is provably convergent under mild conditions. In particular, the proposed method is shown to be convergent with an $O\left(\frac{1}{t^{2}}\right)$ convergence rate in terms of objective function values of the associated convex QP problem, where $t\geq 1$ denotes the iteration counter. Moreover, existing literature also suggests that a restarted scheme could improve the practical performance of the proposed method significantly. Hence, we also analyze the convergence properties of a restarted version of the algorithm and establish an $O\left(\frac{(\log_{2}t)^{2}}{t^{2}}\right)$ convergence rate. Finally, our numerical experiments confirm the efficiency of the proposed algorithm and the scalability of its parallel implementation. Specifically, when applied to strictly diagonal dominant systems, our accelerated algorithm empirically outperforms both the classical Jacobi method and weighted Jacobi method with optimal weight. Furthermore, in numerical tests involving Laplacian systems and general symmetric positive definite systems, our method demonstrates better empirical performance compared to the classical CG method, and shows comparable results as the preconditioned CG method using a simple diagonal preconditioner. Additionally, the speed-up tests for the parallel implementation of our method exhibit promising scalability. These combined theoretical and numerical results provide compelling evidence for the potential applicability of our algorithm in a wide range of real-world applications.
@@ 2 The Nesterov’s accelerated method
In this section, we first revisit the Jacobi-type algorithmic frameworks and show that they are special cases of the general variable-metric gradient methods for solving the associated unconstrained convex QP problem. Hence, the convergence results from the optimization perspectives can be directly applied. Then, we consider the Nesterov’s accelerated gradient method with provable convergence. We also present the convergence results of a restarted version of the former algorithm. We should mention that the analysis for the Nesterov’s accelerated gradient method and its restarted variants is well-developed and mature in the literature; see e.g., [5, 3, 31, 20, 27, 34, 48, 26, 29] and references therein. However, existing analysis is mainly for a general convex optimization problem, whereas the focus of the present paper is an unconstrained convex QP problem. To make the paper self-contained, we also provide a concise analysis here.
@@
where $t$ is the iteration counter. The standard condition based on the fixed-point theory for the Jacobi method to be convergent is $\rho(D^{-1}(L+L^{T}))<1$ where $\rho(\cdot)$ denotes the spectral radius [14]. As a consequence, the Jacobi method is not always convergent for general SPSD systems given an arbitrary initial point, as mentioned before. In order to derive convergent Jacobi-type methods, one may consider a linear combination between the current point $x^{t}$ and the point given by the Jacobi method:
@@
where $\omega\in\left(0,\frac{2}{\lambda_{\textrm{max}}(D^{-1}Q)}\right)$. The above scheme is usually called the weighted Jacobi method, and it is convergent for any given initial point [38] when solving symmetric positive definite systems. We can derive the optimal weight, denoted as $\omega_{\textrm{opt}}$ by minimizing the spectral radius $\rho(I-\omega D^{-1}Q)$ and it is given by
@@
(5) $x^{t+1}=\operatorname{argmin}\left\{f(x^{t})+\left\langle\nabla f(x^{t}),x-x^{t}\right\rangle+\frac{1}{2}\left\lVert x-x^{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}t}}\right\rVert_{S}^{2}\;:\;x\in\mathbb{R}^{n}\right\},\quad\forall\;t\geq 0,$
where $S:=D$ for (3) and $S:=\frac{1}{\omega}D$ for (4), provided that $D\in\mathbb{S}_{++}^{n}$ and $\omega>0$. Here, the term $\frac{1}{2}\left\lVert x-x^{t}\right\rVert_{S}^{2}$ is usually called the proximal term. Note that the traditional gradient method corresponds to the setting $S=\frac{1}{\eta}I_{n}$, where $\eta>0$ is the step size. Therefore, the above updating rule is a special case of the gradient method by changing the standard Euclidean metric to the metric induced by $S$. Hence, these types of methods are also known as preconditioned gradient methods. However, the gradient method specified by (5) is not always convergent. In other words, one needs to choose the matrix $S$ appropriately for convergence. For a given $S\in\mathbb{S}_{++}^{n}$, observe that the updating rule in (5) has the following form:
@@
where $x^{*}$ is the unique solution of (1). The above convergence results coincide with those classical results in the field of optimization; see, e.g., [32, Theorem 2.1.15]. Unfortunately, when the condition number of $Q$, namely $\frac{\lambda_{\textrm{max}}(Q)}{\lambda_{\textrm{min}}(Q)}$, is large, the convergence rate could be rather slow. Moreover, when $Q$ is only SPSD, the above results are no longer applicable. In this case, we can still rely on the convergence results for convex optimization. For instance, suppose that $Q\in\mathbb{S}_{+}^{n}$ and $S=\frac{1}{\eta}I_{n}$, and $\eta=\frac{1}{\lambda_{\textrm{max}}(Q)}$, then
@@
which establishes an $O\left(\frac{1}{t}\right)$ convergence rate in terms of the objective value gap; see [32, Corollary 2.1.2] for more details. Note that the step size $\eta$ depends on $\lambda_{\textrm{max}}(Q)$ which requires additional computational cost to estimate.
@@
The Nesterov’s accelerated framework is recognized to be one of the most influential and efficient first-order methods in the literature. By the description of Algorithm 1, we see that the algorithm is easily applicable and simple to implement. Moreover, there are several major differences between Algorithm 1 and the gradient method (5). First, instead of using a linear approximation, i.e., $f(x^{t})+\left\langle\nabla f(x^{t}),x-x^{t}\right\rangle$, of the objective function $f$ at the current iterate $x^{t}$, we keep using the function $f$ that is in fact an exact second-order approximation. Second, the matrix $S$ is no longer required to be positive definite. Instead, only $S\in\mathbb{S}_{+}^{n}$ is required. Finally, the proximal term is changed to $\frac{1}{2}\left\lVert x-y^{t}\right\rVert_{S}^{2}$ where $y^{t}$ depends on $x^{t-1}$ and $x^{t-2}$ which essentially adds a momentum term to the update of the gradient method (5) to achieve acceleration. The above features are some of the main reasons to achieve accelerated convergence rate, as we will see shortly.
@@ 2.1 Convergence properties
$f(x)-f(x^{t})\geq$ $\;\frac{1}{2}\left\langle x,Qx\right\rangle-\left\langle b,x\right\rangle-\frac{1}{2}\left\langle x^{t},Qx^{t}\right\rangle+\left\langle b,x^{t}\right\rangle-\frac{1}{2}\left\langle x^{t}-y^{t},S(x^{t}-y^{t})\right\rangle$
@@
where we have used the facts that $f(x)-f^{*}\geq 0$ for $x\in\mathbb{R}^{n}$, for $t\geq 1$. Then, by multiplying $\alpha_{t}$ on both side of the last inequality and using the fact that $\alpha_{t-1}^{2}=\alpha_{t}(\alpha_{t}-1)$, we obtain that
@@
$=$ $\;\frac{1}{2}\left\langle\alpha_{t}x^{t}-(\alpha_{t}-1)x^{t-1}-x^{*}{-w^{t-1}},S(\alpha_{t}x^{t}-(\alpha_{t}-1)x^{t-1}-x^{*}{-w^{t-1}})\right\rangle$
$\;{+}\left\langle{w^{t-1}},S({\alpha_{t}x^{t}}-(\alpha_{t}-1)x^{t-1}-x^{*}{-w^{t-1}})\right\rangle$
@@
where in the second equality we used the fact that $\alpha_{t}y^{t}-(\alpha_{t}-1)x^{t-1}-x^{*}=w^{t-1}$, for $t\geq 1$, which can be deduced directly from Line 4 and Line 5 in Algorithm 1, and in the third equality we substituted the definition of $w^{t}:=\alpha_{t}x^{t}-(\alpha_{t}-1)x^{t-1}-x^{*}$, for $t\geq 1$, as given in the statement of the lemma.
@@
$0\leq f(x^{t})-f^{*}\leq\frac{2\left\lVert x^{0}-x^{*}\right\rVert_{S}^{2}}{(t+1)^{2}},\quad t\geq 1.$
@@
We next claim that $\alpha_{t}\geq(t+1)/2$ for all $t\geq 1$. We prove this by mathematical induction. For $t=1$, we see that $\alpha_{1}=(1+\sqrt{1+4\cdot 0})/2=1\geq(1+1)/2$. Hence, the statement is true for $t=1$. Suppose that the statement is true for $t$, i.e., $\alpha_{t}\geq(t+1)/2$. Then, it holds that
$\alpha_{t+1}=\frac{1+\sqrt{1+4\alpha_{t}^{2}}}{2}\geq\frac{1+\sqrt{1+(t+1)^{2}}}{2}>\frac{1+\sqrt{(t+1)^{2}}}{2}=\frac{t+2}{2}.$
Therefore, the statement is true for $t+1$, which implies that the statement is true for all $t\geq 1$. From this, we see that
$0\leq f(x^{t})-f^{*}\leq\frac{\left\langle w^{0},Sw^{0}\right\rangle}{2\alpha^{2}_{t}}\leq\frac{2\left\lVert x^{0}-x^{*}\right\rVert_{S}^{2}}{(t+1)^{2}},\quad t\geq 1.$
Hence, the proof is completed.
Theorem 2.5 gives an improved $O\left(\frac{1}{t^{2}}\right)$ convergence rate that is obviously better than the traditional gradient method, whose worst-case convergence rate is $O\left(\frac{1}{t}\right)$. However, careful readers may observe that the generated objective values $\{f(x^{t})\}$ are not necessarily monotonically decreasing. This is reasonable since Theorem 2.5 only provides a sequence of upper bounds for the objective values. Since the objective function is only assume to be convex, the convergence of the objective values does not imply the convergence of the sequence $\{x^{t}\}$ to an optimal solution. A certain modification to the Nesterov’s accelerated method can be applied to guarantee the convergence of $\{x^{t}\}$; see e.g., [9]. Furthermore, it is worth noticing that, when $Q\in\mathbb{S}_{++}^{n}$, linear convergence can be obtained for a modified Nesterov’s accelerated method [46, Example 3.7].